Jan 10th, Monday
~19:00-22:00; Angel: we took advantage of technical time (T06-21) yielded by Manuel Castillo for NIKA2 general checks and calibrations. Acquisition restarted at ~19:10 after cooling down checks by Dave John and Santiago Navarro. Tau(225 GHz) about 0.25, and quite a stable atmosphere with baseline variations between 1 and 3 Jy/beam (2 mm) during this special shift. After checking the right imbfits file creation, we pointed and focused on Uranus. We got FWHMs 17.2"x17.8" (2 mm); 11.0"x11.6" (1 mm); the antenna seems to be a bit unrelaxed according to the maps. Meanwhile, we checked pointing corrections in the ~150 < Az. < 300 deg and ~20 < El. < 65 deg. ranges. Mean corrections from 11 pointing scans are 6.37" (rms 1.17") in Az. and 2.55" (rms 2.38") in El., and the corrections depend on both coordinates, according to a report of Robert Zylka. For focus we got -0.6 +/- 0.6 mm (stable and close to the expected values after the intervention of PT100 sensors by Salvador Sánchez in the afternoon). Pointing and focus correction on NGC7027. A set of calib_1scan maps were obtained on that source (20220110s181), MWC349 (s182), and CRL2688 (s183). Very stable at the end of the shift. Coming back to Uranus. Resonances were reworked before pointing and a new focus correction. Beam map (h option) on Uranus (s198).
After cooling down, the traces of arr1 showed correlated spikes at ~3, 6, and 9 Hz of unknown origin (a screen-shot below). Dave performed some tests and checked electronics and pre-amplifiers, but without qualitative changes. Possible effects on the maps should be investigated. The behaviour of the other arrays looks fine.
Jan 11th, Tuesday
[AB] A note at the beginning: A new set of corrections to the parameters of the NIKA2 pointing model based on the idem session of early December will be installed by Manuel asap. On the other hand, during the maintenance slot, the telescope engineers have checked in deep the temperature sensors of the antenna back-structure. In the past week, the automatic compensation of temperature gradients was turned off several days due to sensor bad readings and the observers had to increase the frequency of pointing and focus corrections (see the AoD report of Gabriel Paubert). In principle, the system should work appropriately. Otherwise, I will ask you for increasing the frequency of such corrections. We'll be pending on this issue.
16:00-21:00; Angel, Matt: After restart DAQ and re-tune the KIDs, we started observing IMEGIN target NGC628 (160-16) with tau borderline ~0.25. Calib_1 scan of Uranus was completed (168), and returned to IMEGIN, however tau went above criteria of 0.25. Switched to project 137-21 but atmospheric stability declined so waited for 30 mins. The atmosphere became more stable, another calib_1 scan of NGC7027 was taken before starting project 137-21 with tau~0.28.
20:45-03:00; Alex : continuing on 137-21 / WB89-380 - WB89-399 - WB89-437 ; nikaw-21 / @skydip (2022011s203) + focus on 0355+508 ; switching to 192-16 / COSMOS (2 rounds). Time between the end of '@nkpoint f' and the start of PIIC processing: 1min46s, 1min 54, 2min, 4min 16, 31s, 1min14, 5min49.
Jan 12th, Wednesday
03:00 - 05:00; Emmanuel, Juan, Corentin : Tau(225 GHz) between ~0.20 and 0.25 and very stable, during the first hour we did one pointing on 1156+285 and we launched a scan of PSZ2-G228 but an electronic issue came up with the second matrix --> impossible to proceed the observations. Last scans (20220112s25 and s26) likely corrupted.
05:20 - 10:45; Frédéric, Angel: Switching to EMIR backup project 170-21. Targets observed: 38659-low (lines detected after 3 reps.); 54706-low (lines detected after 2 reps.); 53609-low (lines detected, 3 reps.). Arr2 crate issue fixed by Dave John after restarting the corresponding crate by software. Acquisition/imbfits creation restarted after that. Very large (~25"!) pointing corrections after first attempts. Inclinometers updated. Right positioning of the mirrors checked. Wobbler system deactivation verified. Expected pointing corrections recovered. So, scans 20220112s69 to s79 are useless. Karyne Demyk took over at 10:45.
10:45 - 15:00; Karine: Instrument re-tuned at the beginning of the shift. Focus on Mars, a calib_1scan (2022012s86), and a skydip (s87) at the beginning of the shift. Then I switched to project 109-21 for the rest of the shift and did 2 Reps on B68 then 2 Reps on L694. Tau was about 0.11, then up to 0.13-0.15 with some variations.
15:00 - 21:00; Matt: Initial calibrations on 0234+285, tau ~0.125 although light snow outside, and atmospheric stability was good. Moved on to NGC628, first obs tau between 0.10 and 0.16, dropped to 0.04 at the end. Did a second obs of NGC628 where tau increased to ~0.2 by the end but still good stability. Two calib_1 scans were then performed one of CRL618 (20220112s165), and one of Uranus (s168) with tau~0.08. Then switched to GASTON, three observations in total (B211 and B213), excellent atmosphere stability < 0.5 Jy/beam, and tau~0.08.
21:00 - 03:00; Guilaine: Finishing a GASTON map for 30 minutes. Manual re-tuning of NIKA2 KIDs after that. Then a hard sequence of focus/pointing despite the wonderful tau (<0.1): Bad beams (too large, 14-15” at the start at 1.2mm) and pointing corrections too large (up to 10”). It ends up that it was snow on the dish. Converging for 22h30 UT. Then, smooth sequence of calibration on Uranus (calib_1scan, beammap and skypdip: scans 20220112s220 to s222). And 3 repets. on N2CLS COSMOS. I’d suggest to the GASTON team to check their data, at least the last map.
Jan 13th, Thursday
03:00 - 09:00; Miren, Corentin : Tau(225 GHz) very stable ~0.08. Switching to LPSZ 199-16 project. Focus with 1226+023. After several pointings, we did 2 reps on PSZ2-G228 and a skydip (20220113s38). Finally, we did several focuses with 1226+023 and pointing every 2 reps in order to do 4 reps on PSZ2G212 (17/26) and the last 2 reps on PSZ2-G228 (8/8). We wanted to do a calibration one scan with MWC349 but it was too low. Some pointings were a bit awkward (a bit elongated beams).
09:00 - 15:00; Karine: KIDs re-tuned at the beginning of the shift. I started on project 109-21 with 2 Reps on L694. Then on nikaw-21 account, I did a pointing/focus on MCW389 (the one on K3-50A was not good) then a calib_1scan on MCW389, CRL2688, and K3-50A, then a skydip (scans 20220113s108 to s111). I switched back to project 109-21 and did 3 Reps on B68 and 1.5 Rep on L694. Tau was excellent and stable during all the shifts (~0.08-0.1) although the beam size has slowly deteriorated with time. Please note that the calib_1scan on Mars (scan s127) has been done on the 109-21 project account. In addition, I did a couple of focuses trying to get a better beam on the project account. I let you decide if they should all count on the project time! Thanks.
15:00 - 21:00; Matt: Start of the shift tau ~0.11, initial calibrations and then onto project 160-16, starting observation NGC628 for IMEGIN. Then switched to project 110-2 NGC1333-C7 1 rep, followed by 1 rep C2, however noticed atmosphere variations were getting worse (even though tau was good) so went over to Uranus to obtain a calib_1scan (scan number 181), followed by skydip, but the atmosphere still continued to have large 10 Jy/beam variations. We therefore switched to GASTON due to poor sky stability but great tau, repeating B211_213 three times. Atmosphere stabilised (probably due to sunset), variations down at ~2.5 Jy/beam at end of shift. In the middle of the GASTON observations did a calib_1scan on CRL618 (scan id: 200).
21:00 - 03:00; Alexandre: Manual re-tuning of the instrument by Angel, switch to nikaw-21 for a pointing and focus on 0224+671, Observed WB89-391 of project 137-21, switch to nikaw-21, pointing / focus / beammap (scan 233) / calib_1scan (scan 234) on Uranus, switch to 110-21 1 rep on NGC1333-C2 (finished), after waiting more than 8min at tau ~ 0.079 to get the pointing correction... switch to 192-16 COSMOS 3 reps. Time between end of scan and start of PIIC processing for pointing or focus scans : 04:26, 00:56, 00:39 00:59 00:45 00:54 02:08 00:56 00:48 00:44 00:52 00:40 08:32 02:44 05:46 05:41
Jan 14th, Friday
03:00 - 09:00; Laurence: Very smooth shift. Tau225 close to 0.1 all along. A bit of wind bursts (and slightly agitated sky) at the beginning then the wind dropped near 5 UT. Last repetition and pointing for project 192-16 completed at about 2:20 UT. Move on to 199-16 project. Last focus was a couple of hours ago: checking focus: very small drift. Observing PSZ1G226, PSZ2-G211, PSZ2-G212 and PSZ2-G160 in turn. Preparing smooth take over for next observer in the last 30 minutes. Focusing on 1226+023: still very stable despite the Sunrise.
09:00 - 15:00; Karine: I started on project IMEGIN (160-16) with two repetitions on ngc4312 and 2 repetitions on ngc3938. NIKA2 KIDs re-tuned at 10:20. On the nikaw-21 account, I did a pointing/focus/calib_1scan on NGC7027 (scan 100) and a calib_1scan on MWC349 (scan 101), then switched to project 109-21 on B68 (2 repetitions) and did a pointing/focus/calib_1scan on Mars (scan 117) on nikaw-21 account, then back to project 109-21 for 3 repetitions on L694. Tau was very good from 0.08 to 0.15 at the end of the shift. The atmosphere was a bit unstable from noon and the beam got wider.
15:00 - 21:00; Golshan: Opacity slightly below 0.1 along the shift. A suddenly, sometimes crazy, unstable atmosphere before sunset. Three reps. on NGC628 (IMEGIN), the last ones with a scan velocity of 85"/s. Very stable after sunset (2-3 Jy/b at 2 mm). Switch to project 110-21. Three reps. in a row on NGC1333-C2, and 3 reps. on B213-C16. Two calib_1scan on Uranus (scans 161 and 176).
21:00 - 03h00; Nico: Re-tuning at 21h. Started by focus and pointing near G202 under nikaw-21. Switched to 143-21 at 21h30 LT and launched observe_G202. Everything went fine, after a focus, doing a beammap on Uranus (22h45 LT), calib_1scan and skydip (scans 204, 205, and 206, respectively). Now pointing on 0333+321 before observing IC348-1 for project 110-21 (23h34 LT). Pointing corrections are kind of large (6.5, 9.3), it was already about (7,8) when I started the shift. Is everything ok with the pointing model? Smooth obs. of IC348-1. Now switching to COSMOS (01h00 LT).
[AB] Pointing corrections are large indeed, and the reported variations could be attributed to their dependence on both coordinates (see the note of past Mon. 10th about this). However, raw taumeter records show some departures from the mean trend of this part of the night between ~2h and 03h. At the beginning of the next Pool week, updated corrections to the pointing model -based on the pointing session of early December- will be implemented by Manuel Castillo.
Jan 15th, Saturday
03:00 - 09:00; Frédéric: Very smooth shift. Tau225 between 0.1 and 0.2 all along. Last repetitions and pointing for project 192-16 completed before moving to 199-16 project. I have done 1 rep on PSZ1G226 (14/20); 2 reps on PSZ2G211 (14/18); 4 reps on PSZ2G212 (26/26) which is now COMPLETED; 1 rep on PSZ2G160 (12/18). Moving to 160-16 at 8:11 am local. Pointing on 1226+023. Beam was very elliptical. Refocus. Beam ok but a bit large. I have launched 1 repetition on NGC4321.
09:00 - 15:00; Karine: I started on project 160-16 with two repetitions on ngc4312 and 2 repetitions on ngc3938. KIDs re-tuning at 10:30. On the nikaw-21 account, I did a pointing/focus/calib_1scan on Mars then a skydip. On project 109-21, I did 3 repetitions on B68 and 1.5 repetitions on L694. The beam shape has deteriorated over time and several focus calibrations could not really improve it. Tau was very good and stable (0.15) as well as the atmosphere.
15:00 - 21:00; Guilaine: From 15h00 to 17h10, Pointing/focus on several sources. Antenna takes time to relax. Beams are ugly and are improving progressively. 17h11: 64.8 minutes on NGC 628 (IMEGIN 160-16), scanning 60”/sec, excellent conditions (atmosphere very stable). Pointing/focus. Move to project 110-21 and do 2 repetitions (44 minutes) on B213-C16. Pointing/focus/calib_1scan on CRL618 (beams are 11.95”, 17.57”, 11.57”). 1 more repetition (22 minutes) on B213-C16. [AB]: Planned scans for NGC628 were reached during this shift. Please do a check.
21:00 - 03:00; Matt: Start off the evening shift with tau ~0.15, and good atmospheric stability ~0.5 Jy/beam. KIDs retuning at 21:10. The first science target was G202 for project 143-21 for one repetition. We then switched to Uranus for calibrations including pointing, focus, beam map (scan id: 238), calib_1scan (scan id: 239), and skydip (scan id: 240). Then switched to IMEGIN (160-16) to observe NGC3184, after performing initial pointing/focus. Three repeats of NGC3184 were completed, tau stayed constant around 0.16 with good stability ~1 Jy/beam throughout. For the last hour of the shift, we switched to project 192-16. We took initial calibrations and then started observing the COSMOS field with tau ~0.15 and excellent weather stability ~2.0 Jy/beam.
Jan 16th, Sunday
03:00 - 09:00; Juan: Excellent weather tau! 0.15-0.17 very stable, sometimes below 1 Jy/beam fluctuations. KIDs OK no need to re-tune. Observations: COSMOS for 1h30 (3 repetitions of 2 scans in total); PSZ2G211 for 1 hour (3 repetitions of 4 scans), PSZ2G087 (6 repetitions of 4 scans).
09:00 - 15:00; Karine: I started with one repetition on B68 on project 109-21. Angel noticed that the snow cannon was in use right in front of the antenna observing at low elevation... so I switched to nikaw-21 for some calibrations: pointing/focus/calib_1scan on NGC7027, calib_1scan on MWC349 and CRL2688, then a skydip (scans 104 to 107). Then back to project 109-21 for 3.5 repetitions on L694 with many pointing and focus to check the beam shape and size which were better than yesterday. Tau was good and stable (0.15-0.17) as well as the atmosphere.
[AB]: The daytime Operator, Juan Luis, called CETURSA for asking about the snow cannon. They reoriented the cannon but depending on the wind speed/direction, artificial snow streams could affect the quality of the observations.
15:00 - 21:00; Xavier: Opacity excellent during the whole shift. No incident. Very smooth run. After calibration on Neptune and Uranus (focus, calib_1scan: s156 and s164) and one skydip (s157), we started project 110-21 with NGC1333_C7 for 1h30 then did a calibration at CRL618 (s182), and finished today with the GASTON project 122-16 and source B211_213. The beam at 14.4&18.7 arcseconds is not unusual in this period of the day. We hesitated to switch to Emir but decided instead to go on 110-21. The beam improved during that session even during sunset and standard values are recovered on CRL618: 11.9&17.8 arcseconds. KIDs re-tuning at 20:45.
21:00 - 03:00; Matt: Start off the evening shift with tau ~0.14, and good atmospheric stability ~2.0 Jy/beam. First science target was G202 for project 143-21 for one repetition. We then switched to Uranus for calibrations including: pointing, focus, beam map (scan id 212), calib_1scan (scan id: 213), and skydip (scan id: 214). Then switched to IMEGIN (160-16) to observe NGC3184, after performing initial pointing/focus. Three repeats of NGC3184 were completed, tau stayed constant around 0.14 with good stability ~1-2 Jy/beam throughout. For the last hour of the shift, we switched to project 192-16. We took initial calibrations and then started observing the COSMOS field with tau ~0.14 and excellent weather stability ~2.0 Jy/beam.
Jan 17th, Monday
03:00 - 09:00; Emmanuel: Tau about 0.15 over the shift. Stable. Starting with project 192-16: 2 reps. on PSZ2-G226, 1 rep. on PSZ2-G211, and 6 reps. on PSZ2-G133. KIDs re-tuned at 07:45. Switching to project 160-16: 2 reps. on NGC 4321; the last one continued on the next shift.
09:00 - 15:00; Karine: Karine: I started with one repetition on ngc3938 for project 160-16. Then I did a pointing/focus/calib_1scan (scan s101) on Mars on the nikaw-21 project before doing one repetition on B68 for project 109-21. Then I switched project 137-21 for one repetition on IRAS19423+2541. Calibration session on the nikaw-21 account: pointing/focus/calib_1scan on NGC7027 and a calib_1scan on CRL2688, then a skydip (scans s137 to 139, respectively). I ended the shift with one repetition on L694 for project 109-21. Everything was fine: stable atmosphere and good tau, slightly increasing from 0.13 to 0.17.
[AB]: DAQ crashed when scan s115 was running. DAQ restarted quickly at 11:20.
15:00 - 21:00; Longji: I took over the shift at 15:00 with unrelaxed antenna. I kept monitoring the beam and focus for 1.5h (with 1 skydip in the middle, scan number 169) but it didn't improve much so I let Angel switch to EMIR backup project 174-21 and observed target 02524 until sunset. We switched back to NIKA2, made the pointing, beam, and focus again, and found satisfactory results at around 19:00. After making 1 calib_1scan on the secondary calibrator CRL618 (scan number 214), we started observing B211-213 in L_1495 (122-16) for 1 repetition. The shift ended with today's first scan on G202 (143-21). In the middle of the shift, some tracking issue happened on the telescope but it was fixed quickly and didn't significantly affect the observation. [AB]: VAC5 restarted at 18:15.
21:00 - 03:00; Matt: Start off the evening shift with tau ~0.16, and good atmospheric stability ~2.0-3.0 Jy/beam. First science target was G202 for project 143-21 for one repetition. We then switched to Uranus for calibrations including: pointing, focus, beam map (scan id 235), calib_1scan (scan id: 236), and skydip (scan id: 237). During calibrations weather worsened slightly to tau~0.2 (spike to 0.22), with sky stability ~4 Jy/beam, however seemed to settle back to tau~0.16 before next science observations. Switched to IMEGIN (160-16) to observe NGC3184, after performing initial pointing/focus. Over next half-hour tau improved but stability reduced to +/-5 Jy/beam. Four repeats of NGC3184 were completed, tau stayed constant around 0.15, but stability gradually improved to ~2 Jy/beam. For last hour of the shift, we switched to project 192-16. We took initial calibrations and then started observing the COSMOS field with tau ~0.145 and excellent weather stability ~2.0 Jy/beam. KIDs re-tuned at 23:36.
Jan 18th, Tuesday
03:00 - 09:00; Laurence: Smooth shift. Good stability and tau225 ~ 0.15 throughout the shift. Moving on to project 199-16 in order to complete PSZ2G226 while at elevation above 40 deg. Then moving back to project 192-16 on Cosmos field for 2 rounds at elevation between about 50 and 45 degrees. Doing a skydip (scan s46). Switching to project 199-16 to observe PSZ2G133 and PSZ2G160.
09:00 - 10:30: Angel: Targets of LPSZ continued: 2 additional reps. on PSZ2G160. Closing the week with calib_1scans on NGC7027, MC349, and CRL2688 (scans 108, 109, and 110).
Many thanks to the observers who participated in this successful week.